# Review of Machine Learning Approaches for Diagnostics and Prognostics of Industrial Systems Using Industrial Open Source Data

### v2

@@ 4.1 Challenges
In the quest to advance predictive maintenance strategies for ion mill etching systems within semiconductor manufacturing, the diagnosis and prognosis of three principal failure modes—Flowcool leak (F1), Flowcool Pressure Too High Check Flowcool Pump (F2), and Flowcool Pressure Dropped Below Limit (F3)—are critical. Successful prediction of these failures helps in the timely scheduling of maintenance operations, thereby reducing downtimes and enhancing operational efficiency. The 2018 PHM North America data challenge emphasized the analysis of fault behavior within the ion mill etch tool, tasking participants to develop a model from the sensor-derived time series data capable of accurately detecting, diagnosing, and prognosticating the time-to-failure for these malfunctions.
@@
Investigation into the proposed solutions reveals a pronounced inclination towards tree-based methodologies for fault detection and diagnosis. [118, 38, 124] and the team led by [34] have each proposed strategies based on random forest algorithms for early degradation mode detection and diagnosis. Extending the exploration of ML techniques, [93] evaluated an assortment of models, including Logistic Regression, Generalized Linear Models, Multivariate Adaptive Regression Splines, Support Vector Regression (SVR), MLP, random forest, etc. Among these, the random forest model distinguished itself with superior performance. Meanwhile, [126] utilized a knowledge distillation approach towards fault detection across various modes, aiming to bolster the detection of infrequent yet critical faults.
@@
Recent advancements in RUL prediction [48] have predominantly hinged on the application of DL algorithms to analyze complex multivariate time series data. [118, 38, 34] have all utilized Long Short-Term Memory (LSTM) networks and their variations, such as Gated Recurrent Units (GRU) and LSTM-based Metric Regression (LSTM-MR), to capture important features from the raw data. [35] expanded upon this approach by integrating a Temporal Convolutional Network (TCN) with LSTM, enhanced by an attention mechanism, which facilitated refined feature extraction from sensor data for accurate RUL prediction. What’s more, Liu et al.’s two-stage deep transfer learning framework aimed at achieving accurate RUL prediction. In the first stage, the developed model leveraged TCN for initial temporal feature learning, followed by domain adversarial learning for data alignment based on one fault mode. Then in the second stage, the first-stage model was fine-tuned based on other fault mode data to handle multiple fault modes and enhance the RUL prediction performance [66]. Distinct from these methods, [124] explored transformer networks, focusing solely on data from abnormal operation phases, while [68] provided a comprehensive comparison of state-of-the-art methods, including TCN, LSTM, attention-based mechanisms, CNN, and Transformer, culminating in the development of the CeRULEo library to facilitate research in the field. Moreover, [93] introduced a novel approach using DTW for RUL estimation, leveraging a library of truncated degradation curves and health score models to refine the final RUL predictions. Such diversity in methodologies not only underscores the complexity of fault diagnosis and prognosis in high-tech manufacturing but also showcases the robustness of ML in tackling these challenges.
4.3 2019 PHM NA (Fatigue Crack)
@@
The 2019 PHM Conference Data Challenge put a spotlight on the task of fatigue crack length estimation and prediction within aluminum structures at different points. The challenge harnessed wave signal data collected by piezoelectric sensors subjected to both static and dynamic tensile stresses. Given that the test dataset was limited to wave signals from several initial loading cycles, participants were challenged to estimate crack lengths where signal data were present and predict future crack growth for specified cycles lacking signal data. The complexity of the problem delineated a clear path: data-driven methodologies were viable when signals existed, whereas scenarios lacking wave data required exploration into physics-based strategies. Reflecting this path, solutions derived from this challenge have predominantly employed a hybrid approach of data-driven and physics-based techniques.
@@
Regarding the estimation of fatigue crack length with accessible wave signals, a variety of data-driven models and feature engineering strategies emerged, both within and beyond the competition’s scope. [46] proposed a neural network architecture reliant on features manually derived from raw signals, such as the Pearson correlation coefficient, phase shifts, energy, and information entropy, to train models for accurate crack length estimation. Similarly, [50] initiated their approach with signal preprocessing, utilizing techniques like band-pass filtering and phase alignment to mitigate noise and uncertainty before applying physically insightful feature extraction methods. Thereafter, they harnessed a random forest algorithm, optimizing it through feature selection and grid search for hyperparameter fine-tuning, to estimate crack lengths. Consistent with these methods, [121] also leveraged a band-pass filter for feature extraction from raw wave signals, subsequently constructing an SVR model with hyperparameters optimized via grid search. Rao and collaborators, meanwhile, designed an ensemble learning regression model to improve estimation performance with four useful extracted features (root mean square value, correlation coefficient, first peak value, and the logarithm of kurtosis) [86].
@@
4.4 2020 PHM EU (Filtration System)
In the domain of industrial maintenance, filtration systems are pivotal for helping process pollutants from industrial equipment, ensuring seamless system operation. A predominant challenge encountered within these systems is filter clogging - a phenomenon where accumulated pollutants impede flow rates, thereby disrupting standard industrial workflows. Addressing this issue, the PHM-Europe Conference 2020 data challenge concentrated on the prediction of filtration systems’ RUL, crucial for timely maintenance and replacement scheduling. RUL, in this context, is delineated as the time until the pressure differential across the filter breaches a threshold of 20 psi. The challenge provided the PHME20 dataset, collected from a controlled experimental setup designed to simulate filter clogging at various contamination levels. Twelve distinct conditions were established influenced by two operational parameters: solid ratio (%) and particle size (µm).
The champion solution [67] employed a novel hybrid approach, combining kernel regression with fundamental statistical methodologies. This approach, underscored by complex data analysis, proved efficacious in enhancing the RUL prediction. Meanwhile, the runner-up, [10], adopted different cutting-edge feature engineering, and ML methods. Their process began with the extraction of features through a rolling window technique and proceeded with feature selection informed by the linear kernel Support Vector Machine (SVM) coefficients, RFE, correlation matrix, and monotonicity testing. A four-layer sequential neural network was their model of choice, supported by K-fold cross-validation throughout the training and validation stages. Capturing the third position, [43] conducted a comparative study of tree-based algorithms and the Bayesian approach. Specifically, random forest, gradient boosting, and Gaussian process regression were utilized to estimate the RUL of the filtration system, supplemented by a novel fault-based RUL assignment that integrated ”Piecewise RUL Assignment” and ”Linear RUL Assignment”.
@@
The 2021 North America PHM Data Challenge was primarily focused on the prediction of RUL for turbofan engines [13], specifically under four distinct flight conditions and seven failure modes. Participants were required to create predictive models leveraging the N-CMAPSS dataset, aiming to accurately forecast RUL using complex condition monitoring data. The dataset, a collection consisting of 90 synthetic run-to-failure trajectories for training and an additional 38 truncated datasets for testing, served as a comprehensive foundation for developing robust predictive algorithms to predict RUL accurately.
@@
During the competition, the winning team, led by Lovberg, put forward an innovative approach leveraging a deep convolutional neural network (DCNN) [69]. This network was distinguished by its use of dilated convolutions complemented by gated linear unit activations and integration of residual skip connections. Those techniques were designed to expand the network’s receptive field and enhance flexibility, so as to reduce the complexity of the neural network architecture by using less number of parameters, but still having comparable performance. Additionally, they adopted a strategic sequence sampling method, minimizing less informative samples while retaining enough degradation signals for the network’s input. This methodology distinguished their solution and got 1st place in the PHM data challenge competition. Moreover, Solis-Martin et al. and DeVol’s team pursued advancements in DCNNs as well. The former developed a two-level DCNN system where the first-level DCNN focused on extracting vital features from raw data and the second-level DCNN leverages the output from the first level to accurately estimate the RUL [97]. The latter drew upon established DL architectures, deploying the basic blocks or modules from the VGG, GoogLeNet, and ResNet designs into their DCNN framework. This enabled a comparative analysis of model performances using a variety of well-known architectural features [24, 25].
@@
Outside of the competition, [19] addressed concerns regarding the uncertain and poor interpretability of deep learning models by integrating principal component analysis (PCA) to refine time-domain feature sets and subsequently applying four supervised learning techniques, including artificial neural network, random forest, extreme random forest and XGBoost, to estimate RUL. Their innovative use of a custom loss function in conjunction with traditional ANN models got the best results in both Area Under the Receiver Operating Characteristic (AUROC) and Area Under the Precision-Recall (AUPR) metrics. Moreover, the domain of PHM has witnessed an upsurge in the application of XAI techniques to bolster the interpretability and trustworthiness of ML models. Various methods, including LIME, SHAP, LRP, Image-Specific Class Saliency Maps, and Gradient weighted Class Activation Mapping (Grad-CAM), have been reported in recent literature [18, 20, 98]. These techniques are increasingly critical in elucidating the decision-making processes of complex models, thereby fostering greater confidence in their predictive capabilities.
@@ 5 Challenges and Possible Solutions
By reviewing nine open PHM data challenge competitions over the past six years, these competitions have showcased a variety of ML approaches applied to different industrial systems. In this section, we summarize common challenges regarding data-related issues and model-related issues and analyze relevant solutions applied to solve these challenges. Moreover, we evaluate the limitations of these competitions and suggest future directions for their evolution, aiming to accelerate the development of next-generation ML-driven PHM solutions.
@@
Data-related challenges encompass issues such as missing data, data imbalance, and domain shift while model-related challenges include the critical aspects of model selection, interpretability of ML models, and their robustness and generalization capabilities. These challenges highlight the complexities of developing effective ML methods to solve various PHM problems. The limitations of current PHM competitions reveal a gap in adopting systematic approaches for building effective PHM systems, and a need for multi-modal machine learning analysis.
5.1 Data-related Issues
5.1.1 Missing Data
Handling missing data in ML, especially in the context of the PHM is significant, as incomplete data can significantly impact the model performance and prediction accuracy. The strategies to address missing data in PHM have evolved, encompassing a range of techniques from basic deletion to advanced imputation methods. This was evident in the 2021/2022 PHM EU competitions, where missing data posed significant challenges. In the fuse test bench dataset (2021 PHM EU), about 10% of the data was missing, and it was not evenly distributed across variables. Meanwhile, the SPI dataset(2022 PHM EU) had missing information in specific fields. To tackle these issues, various approaches were employed.
A straightforward strategy is the deletion of data points with missing values, such as listwise or pairwise deletion. For example, in the 2022 SPI dataset, instances having missing values in crucial identifiers like ”Panel_ID”, ”Figure_ID”, and ”Component_ID” were eliminated [102]. However, this approach can lead to the loss of valuable information to some degree. On the imputation front, several innovative methods were used in the 2021 PHM EU competition. One team applied the Last Observation Carried Forward (LOCF) method, coupled with backward filling, to address the gaps [22]. Interpolation was another technique used to estimate missing values, ensuring that cases with absent data did not skew the results. Additionally, Ramezani’s team introduced a novel approach, PLS-MV, a partial least squares-based method for imputing missing values [85]. In addition to various imputation methods highlighted in the competitions, there remains scope for exploration in future research such as mean/median/multiple, K-Nearest Neighbors(KNN), regression imputation, maximum likelihood estimation, and ML-based approaches [39]. These methods offer a promising avenue for further enhancing data quality and model reliability in PHM systems.
5.1.2 Data Imbalance
Data imbalance arises when the distribution of classes in a dataset is uneven. This issue is especially pronounced in PHM due to the rarity of failure events in comparison to data representing normal operations. Data imbalance can significantly undermine the performance of ML models, particularly in classification tasks. When trained on imbalanced data, models may become biased towards the majority class (normal operation) and may not effectively recognize the minority class (failure). This leads to poor performance in predicting failures, which may bring ineffective maintenance planning and unexpected downtimes. In PHM data challenge competitions, this issue is commonly encountered. Fuse test bench dataset from the 2021 PHM EU competition and printed circuit board dataset from the 2022 PHM EU competition exemplify this issue. The former displayed an obvious imbalance with only 3-4 experiments representing each fault class against 70 experiments for the healthy class. Similarly, in the printed circuit board dataset, 95% of the SPI data was classified as healthy, highlighting a significant imbalance. Additionally, the 2018 PHM NA competition also faced this challenge in the Ion Mill Etching (IME) dataset, where the faulty data is much less than normal operation data.
To counteract data imbalance in PHM, various strategies have been adopted. Resampling techniques are commonly used in competitions to adjust the dataset to balance the class distribution, either by oversampling the minority class or downsampling the majority class. Synthetic data generation is another approach, as demonstrated by a team in the 2021 PHM EU competition that used the SMOTE to create synthetic samples of the minority class to balance the dataset [22]. Advanced algorithms also play a crucial role in addressing data imbalance. For instance, Tang’s team in the 2022 PHM EU competition used the FIR to control the imbalance ratio in each mini-batch during neural network training [104]. Liu et al. employed transfer learning and domain adaptation techniques in the 2018 PHM NA competition to facilitate knowledge transfer across different fault modes, addressing the issue of insufficient data in specific faults [66]. Moreover, during the 2018 PHM NA competition, many teams utilized Random Forest, an ensemble learning method known for its proficiency in handling imbalanced data, by constructing a forest of decision trees. These approaches collectively demonstrate the diverse and innovative methods employed to mitigate the challenges posed by data imbalance in the PHM field.
Beyond the techniques showcased in the competitions, new DL approaches are emerging to address this issue, such as the semi-supervised information maximizing generative adversarial network [117], the integration of deep residual networks with auxiliary classifier generative adversarial networks [14], the combination of DL with SMOTE [21], etc. Furthermore, Aguiar et al. introduced a standardized experimental framework to evaluate 24 state-of-the-art data stream algorithms across 515 imbalanced data streams [1]. Going forward, there is significant potential for the exploration of additional DL-based strategies to enhance the handling of data imbalance in the PHM domain.
5.1.3 Domain Shift
Domain shift refers to the changes in the data distribution between the training phase (source domain) and the real-world application phase (target domain) of the ML models. This phenomenon frequently occurs when models, initially trained on data from a specific set of machines or under certain conditions, are subsequently applied to different machines or varied operating conditions. Such a shift can markedly affect the performance and reliability of ML models in PHM, making tackling with domain shift problem essential for maintaining the robustness and effectiveness of PHM systems. Transfer learning has emerged as a primary solution to domain shift challenges, as it can achieve the adaptation of models from one domain to be effective in another through fine-tuning and domain adaptation techniques. Additionally, robust modeling approaches and the ability to continuously update models with new data are also useful and imperative for mitigating the impact of domain shift.
A clear example of domain shift was observed in the rock drill dataset from the 2022 PHM NA competition. The dataset encompassed training data from five rock drill machines, validation data from another machine, and test data from two additional machines, highlighting potential domain shift issues between training, validation, and test data. Similarly, the 2020 PHM EU dataset demonstrated domain shift, with training data featuring small and large particle sizes, while the test data included medium particle sizes. To address these challenges, various strategies were employed. In the 2022 PHM NA competition, Kim et al. utilized domain adversarial neural networks (DANN) along with the minimization of MMD to tackle domain discrepancies [49]. Meanwhile, Oh et al. employed the deep CORAL method, calculating coral loss to diminish domain discrepancy effects by aligning the covariance of the six training domains with that of the test domain [80]. For the 2020 PHM EU competition, Tian’s team developed a novel TEL framework, facilitating knowledge transfer from the source to the target domain [107]. These approaches illustrate the potential to overcome domain shift problems in PHM, ensuring that models can maintain accurate predictions despite changes in their working or operation conditions.
Table 8: Summary of Methodologies Solving Data-related Issues in PHM Data Challenge Competitions
Data Issue Potential Solution
@@ 5.1.3 Domain Shift
Data Imbalance

### v3

@@ 3.2 Assessment & Prognosis
The 2018 PHM NA emphasized the analysis of fault behavior within the ion mill etch tool, tasking participants to develop a model from the sensor-derived time series data capable of accurately detecting, diagnosing, and prognosticating the time-to-failure for three principal failure modes—Flowcool leak (F1), Flowcool Pressure Too High Check Flowcool Pump (F2), and Flowcool Pressure Dropped Below Limit (F3). Additionally, the Ion Mill Etching (IME) dataset faced an imbalance issue, where the faulty data is much less than normal operation data.
@@
For this dataset, [125, 42, 129, 38]have each proposed strategies based on random forest algorithms for early degradation mode detection and diagnosis. Extending the exploration of ML techniques, [99] evaluated an assortment of models, including Generalized Linear Models, MLP, Multivariate Adaptive Regression Splines, Support Vector Regression (SVR), random forest, etc. Among these, the random forest model distinguished itself with superior performance. Meanwhile, [131] utilized a knowledge distillation approach towards fault detection across various modes, aiming to improve detection performance of infrequent but critical faults.
@@
Recent advancements in RUL prediction [52] have predominantly hinged on the application of DL algorithms to analyze complex multivariate time series data. [125, 42, 38] have all utilized Long Short-Term Memory (LSTM) networks and their variations, such as Gated Recurrent Units (GRU) and LSTM-based Metric Regression (LSTM-MR), to capture important features from the raw data. [39] expanded upon this approach by integrating a Temporal Convolutional Network (TCN) with LSTM with attention mechanisms, which facilitated refined feature extraction from sensor data for accurate RUL prediction. Moreover, Liu et al.’s two-stage deep transfer learning framework aimed at achieving accurate RUL prediction. In the first stage, the developed model leveraged TCN for initial temporal feature learning, followed by domain adversarial learning for data alignment based on one fault mode. Then in the second stage, the first-stage model was fine-tuned based on other fault mode data to handle multiple fault modes and enhance the RUL prediction performance [70]. Distinct from these methods, [129] explored transformer networks, focusing solely on data from abnormal operation phases, while [72] provided a comprehensive comparison of state-of-the-art methods, including TCN, LSTM, attention-based mechanisms, CNN, and Transformer. Moreover, [99] introduced a novel approach using DTW for RUL estimation, leveraging a library of truncated degradation curves and health score models to refine the final RUL predictions.
3.2.2 2019 PHM NA (Fatigue Crack)
@@
2019 PHM NA focused on the task of fatigue crack length estimation and prediction within aluminum structures at different points. The challenge harnessed wave signal data collected by piezoelectric sensors subjected to both static and dynamic tensile stresses. Given that the test dataset was limited to wave signals from several initial loading cycles, participants were challenged to estimate crack lengths where signal data were present and predict future crack growth for specified cycles lacking signal data. The complexity of the problem is that data-driven methodologies were viable when signals existed, whereas scenarios lacking wave data required exploration into physics-based strategies. Therefore, solutions derived from this challenge have predominantly employed a hybrid approach of data-driven and physics-based techniques.
@@
Regarding the estimation of fatigue crack length with accessible wave signals, a variety of data-driven models and feature engineering strategies emerged, both within and beyond the competition’s scope. [50] proposed a neural network architecture reliant on features manually derived from raw signals, such as the Pearson correlation coefficient, phase shifts, energy, and information entropy, to train models for accurate crack length estimation. Similarly, [54] initiated their approach with signal preprocessing, utilizing techniques like band-pass filtering and phase alignment to mitigate noise and uncertainty before applying physically insightful feature extraction methods. Thereafter, they harnessed a random forest algorithm, optimizing it through feature selection and grid search for hyperparameter fine-tuning, to estimate crack lengths. Additionally, [127] also leveraged a band-pass filter for feature extraction from raw wave signals, subsequently constructing an SVR model with hyperparameters optimized via grid search. Rao and collaborators, meanwhile, designed an ensemble learning regression model to improve estimation performance with four useful extracted features (root mean square value, correlation coefficient, first peak value, and the logarithm of kurtosis) [92].
3.2.3 2020 PHM EU (Filtration System)
@@
In the domain of industrial maintenance, filtration systems are crucial for helping process pollutants from industrial equipment, ensuring seamless system operation. A predominant challenge encountered within these systems is filter clogging - a phenomenon where accumulated pollutants impede flow rates, thereby disrupting standard industrial workflows. To address this issue, 2020 PHM EU concentrated on the prediction of filtration systems’ RUL. RUL, in this context, is delineated as the time until the pressure differential across the filter breaches a threshold of 20 psi. The challenge provided the PHME20 dataset, collected from a controlled experimental setup designed to simulate filter clogging at various contamination levels. Twelve distinct conditions were established influenced by two operational parameters: solid ratio (%) and particle size (µm). Additionally, the 2020 PHM EU dataset demonstrated domain shift problem, with training data featuring small and large particle sizes, while the test data included medium particle sizes.
The champion solution [71] employed a novel hybrid approach, combining kernel regression with fundamental statistical methodologies. Meanwhile, the runner-up, [10], adopted different cutting-edge feature engineering, and ML methods. Their process began with the extraction of features through a rolling window technique and proceeded with feature selection informed by the linear kernel Support Vector Machine (SVM) coefficients, RFE, correlation matrix, and monotonicity testing. A four-layer sequential neural network was their model of choice, supported by K-fold cross-validation throughout the training and validation stages. Capturing the third position, [47] conducted a comparative study of tree-based algorithms and the Bayesian approach. Specifically, random forest, gradient boosting, and Gaussian process regression were utilized to estimate the RUL of the filtration system, supplemented by a novel fault-based RUL assignment that integrated ”Piecewise RUL Assignment” and ”Linear RUL Assignment”.
@@
2021 PHM NA was primarily focused on the prediction of RUL for turbofan engines [13], specifically under four distinct flight conditions and seven failure modes. Participants were required to create predictive models leveraging the N-CMAPSS dataset, aiming to accurately forecast RUL using complex condition monitoring data. The dataset, a collection consisting of 90 synthetic run-to-failure trajectories for training and an additional 38 truncated datasets for testing, served as a comprehensive foundation for developing robust predictive algorithms to predict RUL accurately.
@@
During the competition, the winning team, led by Lovberg, put forward an innovative approach leveraging a deep convolutional neural network (DCNN) [73]. This network was distinguished by its use of dilated convolutions complemented by gated linear unit activations and integration of residual skip connections. Those techniques were designed to expand the network’s receptive field and enhance flexibility, so as to reduce the complexity of the neural network architecture by using less number of parameters, but still having comparable performance. Additionally, they adopted a strategic sequence sampling method, minimizing less informative samples while retaining enough degradation signals for the network’s input. Moreover, Solis-Martin et al. and DeVol’s team pursued advancements in DCNNs as well. The former developed a two-level DCNN system where the first-level DCNN focused on extracting important features from raw data and the second-level DCNN leverages the output from the first level to accurately estimate the RUL [103]. The latter drew upon established DL architectures, deploying the basic blocks or modules from the VGG, GoogLeNet, and ResNet designs into their DCNN framework. This enabled a comparative analysis of model performances using a variety of well-known architectural features [25, 26].
@@
Outside of the competition, [19] addressed concerns regarding the uncertain and poor interpretability of deep learning models by integrating principal component analysis (PCA) to refine time-domain feature sets and subsequently applying four supervised learning techniques, including artificial neural network, random forest, extreme random forest and XGBoost, to estimate RUL. Their innovative use of a custom loss function in conjunction with traditional ANN models got the best results in both Area Under the Receiver Operating Characteristic (AUROC) and Area Under the Precision-Recall (AUPR) metrics. Moreover, the domain of PHM has witnessed an upsurge in the application of XAI techniques to enhance the interpretability and trustworthiness of ML models. Various methods, including LIME, SHAP, LRP, Image-Specific Class Saliency Maps, and Gradient weighted Class Activation Mapping (Grad-CAM), have been reported in recent literature [18, 20, 104]. These techniques are useful in elucidating the decision-making processes of complex models.
3.3 Comprehensive Summarization of Data-Driven Approaches in Recent PHM Data Challenge Competitions
After conducting an in-depth review of data competitions from the last six years, Figure 2 provides details into the publication count and the frequency of particular ML approaches within each competition. Furthermore, Figure 3 illustrates the density of specific ML or DL approaches discussed in this paper. It is important to note that we count the occurrences of distinct ML methods mentioned in a research paper. Given that some articles employ multiple ML methods, the aggregate count of methods exceeds the total number of publications.
Data Collection: Data collection is the foundation of these open-source data challenge competitions (PHM Society and IEEE Reliability Society). Companies across various industries contribute datasets to encourage PHM community to develop innovative solutions to address real-world challenges. Typically, these industrial datasets encompass a diverse range of data types, including but not limited to time-series data, tabular data, images, sensor readings, and simulation data.
Data Processing: An important stage before ML model development is data processing, which is crucial for enhancing data quality and ensuring effective model training [17, 111, 21, 35]. This stage can be further subdivided into two aspects: data preprocessing and feature engineering. Data preprocessing addresses raw data challenges, typically including missing data, noise, outliers, data imbalance, and scaling issues. Techniques like imputation [29], denoising [23], outlier detection [77], resampling [16], and normalization & standardization [58, 34] are commonly applied. Feature engineering follows, refining data representation post-preprocessing. This step often uncovers hidden insights and deepens understanding of data, therefore significantly enhancing model performance and predictive capabilities in the later stage [98].
Conventional Machine Learning & Deep Learning Techniques: Post data processing, various conventional ML, and DL methods are developed to solve the competition problems using useful and usable data, as shown in Figure 4. This segment includes ”model training and testing” and ”model prediction and classification”. During the training and testing phase. ML/DL techniques are diverse, including Tree-Based Methods, CNNs, RNNs and their variants, Transformers, DANN, SVMs, Unsupervised Learning, Ensemble Learning, Similarity-Based Methods, Rule-Based Methods, Transfer Learning, Domain Adaptation, etc. [96, 89] Once the ML models are trained and tested well, they could be deployed into various PHM applications which include detecting abnormal performance or faults, classifying different faults, assessing the current health state of systems, or predicting the RUL of components.
@@ 4 Challenges and Possible Solutions
In this section, we summarize common challenges regarding data-related issues and model-related issues and analyze relevant solutions applied to solve these challenges, shown in Table 7. Data-related challenges encompass issues such as missing data, data imbalance, and domain shift while model-related challenges include the critical aspects of model selection, interpretability of ML models, and their robustness and generalization capabilities. These challenges highlight the complexities of developing effective ML methods to solve various PHM problems.
Moreover, the limitations of current PHM competitions reveal a gap in adopting systematic approaches for building effective PHM systems, and a need for multi-modal machine learning analysis. We also suggest further steps, aiming to accelerate the development of next-generation ML-driven PHM solutions.
Table 7: Summary of Common Challenges Regrading Data-related and Model-related Issues
Common Challenges Potential Solutions Related Competitions
@@
2021 PHM EU
2022 PHM EU
Data Issue Data Imbalance
@@
2020 PHM EU
2022 PHM NA
Model Selection
(1) Need to Consider Volume and Quality
of Data
(2) A Trade-off Between Computational
Cost and Performance
All Competitions
Model Issue Model Interpretability
(1) Explainable Artificial Intelligence
(XAI) methods like LIME, SHAP, etc.
No Competition
Required
Model Robustness
& Generalization
(1) Data Augmentation
(2) Regularization Techniques
(3) Ensemble Methods
(4) Transfer Learning & Domain Adaptation
All Competitions
4.1 Data-related Issues
4.1.1 Missing Data
Handling missing data in ML, especially in the context of the PHM is necessary, as incomplete data can significantly impact the model performance and prediction accuracy. The strategies to address missing data in PHM have evolved, encompassing a range of techniques from basic deletion to advanced imputation methods. A straightforward strategy is the deletion of data points with missing values, such as listwise or pairwise deletion. For example, in the 2022 SPI dataset, instances having missing values in crucial identifiers like ”Panel_ID”, ”Figure_ID”, and ”Component_ID” were eliminated [108]. However, this approach can lead to the loss of valuable information to some degree. Regarding imputation methods, several innovative methods were used in 2021 PHM EU. [23] applied the Last Observation Carried Forward (LOCF) method, coupled with backward filling, to address the gaps, while  [91] introduced a novel approach, PLS-MV, a partial least squares-based method for imputing missing values. Interpolation was another technique used to estimate missing values, ensuring that cases with absent data did not skew the results. In addition to various imputation methods highlighted in the competitions, there remains scope for exploration in future research such as mean/median/multiple, K-Nearest Neighbors(KNN), regression imputation, maximum likelihood estimation, and ML-based approaches [43].
4.1.2 Data Imbalance
Data imbalance arises when the distribution of classes in a dataset is uneven. This issue is evident in PHM due to the rarity of failure events in comparison to data representing normal conditions. Data imbalance can significantly undermine the performance of ML models, particularly in classification tasks. When trained on imbalanced data, models may become biased towards the majority class (normal operation) and may not effectively recognize the minority class (failure). This leads to poor performance in predicting failures, which may bring ineffective maintenance planning and unexpected downtimes.
To counteract data imbalance in PHM, various strategies have been adopted. Resampling techniques are commonly used to adjust the dataset to balance the class distribution, either by oversampling the minority class or downsampling the majority class. Synthetic data generation is another approach, as demonstrated by a team in 2021 PHM EU that used the SMOTE to create synthetic samples of the minority class to balance the dataset [23]. Advanced algorithms also play a crucial role in addressing data imbalance. For instance, Tang’s team in 2022 PHM EU used the FIR to control the imbalance ratio in each mini-batch during neural network training [110]. Liu et al. employed transfer learning and domain adaptation techniques in 2018 PHM NA to facilitate knowledge transfer across different fault modes, addressing the issue of insufficient data in specific faults [70]. Moreover, during the 2018 PHM NA competition, many teams utilized Random Forest, an ensemble learning method known for its proficiency in handling imbalanced data, by constructing a forest of decision trees.
Beyond the techniques showcased in the competitions, new DL approaches are proposed, such as the semi-supervised information maximizing generative adversarial network [124], the integration of deep residual networks with auxiliary classifier generative adversarial networks [14], the combination of DL with SMOTE [22], etc. Furthermore, a standardized experimental framework is proposed by Aguiar’s team in order to evaluate 24 state-of-the-art data stream algorithms across 515 imbalanced data streams  [1]. Going forward, there is a need for the exploration of additional DL-based strategies to enhance the handling of data imbalance in the PHM domain.
4.1.3 Domain Shift
Domain shift refers to the changes in the data distribution between the training phase (source domain) and the real-world application phase (target domain) of the ML models. This phenomenon frequently occurs when models, initially trained on data from a specific set of machines or under certain conditions, are subsequently applied to different machines or varied operating conditions. Such a shift can markedly affect the performance and reliability of ML models in PHM, making tackling with domain shift problem essential for maintaining the robustness and effectiveness of PHM systems. Transfer learning has emerged as a primary solution to domain shift challenges, as it can achieve the adaptation of models from one domain to be effective in another through fine-tuning and domain adaptation techniques. For instance, Kim et al. utilized domain adversarial neural networks (DANN) along with the minimization of MMD to tackle domain discrepancies in 2022 PHM NA [53]. Meanwhile, Oh et al. employed the deep CORAL method, calculating coral loss to diminish domain discrepancy effects by aligning the covariance of the six training domains with that of the test domain [86]. Moreover, Tian’s team developed a novel TEL framework, facilitating knowledge transfer from the source to the target domain in 2020 PHM EU [113]. Additionally, robust modeling approaches and the ability to continuously update models with new data are also useful and imperative for mitigating the impact of domain shift.